### v4

# Digital-physical testbed for ship autonomy studies in the Marine Cybernetics Laboratory basin

Abstract
The algorithms developed for Maritime Autonomous Surface Ships (MASS) are often challenging to test on actual vessels due to high operational costs and safety considerations. Simulations offer a cost-effective alternative and eliminate risks, but they may not accurately represent real-world dynamics for the given tasks. Utilizing small-scale model ships and robotic vessels in conjunction with a laboratory basin provides an accessible testing environment for the early stages of validation processes. However, designing and developing a model vessel for a single test can be costly and cumbersome, and researchers often lack access to such infrastructure. To address these challenges and enable streamlined testing, we have developed an in-house testbed that facilitates the development, testing, verification, and validation of MASS algorithms in a digital-physical laboratory. This infrastructure includes a set of small-scale model vessels, a simulation environment for each vessel, a comprehensive testbed environment, and a digital twin in Unity. With this, we aim to establish a full design and verification pipeline that starts with high-fidelity simulation models of each model vessel, to the model-scale testing in the laboratory basin, allowing possibilities for moving towards semi-fullscale validation with R/V milliAmpere1 and full-scale validation with R/V Gunnerus. In this work, we present our progress on the development of this testbed environment and its components, demonstrating its effectiveness in enabling ship guidance, navigation, and control (GNC), including autonomy.
keywords:
Maritime Autonomous Surface Ships (MASS), Simulation, Testing, Validation, Model Testing, Digital Twin, Maritime Digitalization
articletype: ARTICLE
1 Introduction
Autonomous vessels have the potential to significantly influence the maritime sector by enhancing safety, operational efficiency, and economic viability [10, 21]. Recent advances in sensor technology, artificial intelligence, and marine cybernetics have spurred the development of prototype autonomous vessels of various sizes [29]. Additionally, large-scale initiatives have emerged to demonstrate the real-world feasibility of uncrewed ships under diverse conditions [28]. These efforts are largely driven by benefits such as increased safety (through reduced human error), lower operating costs, and reduced emissions, all of which hold significant promise for reshaping shipping operations and logistics [10, 27].
To bridge the gap between concept and full-scale deployment, researchers and industry leaders have developed or repurposed vessels of varying sizes as experimental platforms for autonomous navigation [29]. The milliAmpere1 [13] and milliAmpere2 [5] ferries—both built at half-scale for urban waterways—have facilitated the development and validation of key autonomy functions such as sensor fusion, collision avoidance, and automated docking under real harbor conditions [18]. Yara Birkeland [34] is a fully electric container ship providing a full-scale test platform for uncrewed cargo operations. Furthermore, the crewed research vessel R/V Gunnerus [17] is equipped with conventional ship systems and an extensive suite of sensors and data acquisition tools, and it provides a platform for advanced scientific studies. In the commercial scale, companies like Maritime Robotics and SeaRobotics each offer a fleet of autonomous surface vessels (ASVs) of multiple sizes for robotics applications and hydrographic surveys, while Reach Subsea, Ocean Infinity, and Fugro each develop their fleets of autonomous vessels specifically designed for offshore operations.
To balance realism with efficiency, researchers often adopt a hybrid approach that combines physical trials with software simulations. Test basins with programmable wave makers offer realistic sea conditions for evaluating control systems, but access to such facilities is limited, and building model vessels can be costly and time-consuming. In response to this need, we have developed an infrastructure that supports the design, verification, and validation of algorithms for Maritime Autonomous Surface Ships (MASS) with a growing fleet of Cybership (C/S) vessels. This testbed is intended for use by researchers, broader academia, and industry collaborators. It includes high-fidelity simulations based on hydrodynamic analyses, streamlined simulations for interface testing and early-stage development, and publicly accessible hydrodynamic data to encourage wider adoption and collaboration. Our goal is to provide a comprehensive testbed and a framework to support other laboratories pursuing similar designs.
Figure 1: Overview of Marine Cybernetics Laboratory components.
In this work, we introduce a digital-physical testbed for Maritime Autonomous Surface Ships (MASS) that combines high-fidelity simulations, reduced-order models, and real-time sensor feedback within a unified framework. Our testbed facilitates a smooth transition from simulation to physical experimentation, supported by a digital twin environment in Unity for visualization and remote control. This setup aims to enhance the development and validation of autonomy algorithms by providing a cost-effective and flexible platform. The primary contribution is an open, modular software-hardware stack—featuring curated datasets and a hydrodynamic model library—that transforms the Marine Cybernetics Laboratory basin into a benchmark platform for ship autonomy research.
The paper is organized as follows. Section 2 provides an overview of the testbed infrastructure, including the physical and digital components. Section 3 describes the simulation environments used for algorithm development and testing, including high-fidelity simulations, reduced-order models, and sensor simulations. Section 4 details the physical infrastructure of the MC-Lab, including the wave basin and hardware architecture of the C/S fleet. Finally, Section 5 concludes the paper and outlines future work.
2 Infrastructure
2.1 Overview
The multi-layered infrastructure integrates digital simulations with physical testing environments to create a seamless, progressive workflow for developing and validating MASS algorithms. At the core of this setup is the C/S Fleet (see Table 1), originally developed for education and research projects at NTNU. These vessels form a key component of the laboratory testbed, serving as the primary physical platforms for real-world experimentation.
Vessel Appears on Status
Cybership I [24] End of life
Cybership II [32] End of life
Cybership III [35] End of life
C/S Enterprise I [31] Operational
C/S Saucer [15] Operational
C/S Inocean Cat I Drillship [1] Operational
C/S Voyager [6] [^1] C/S Jonny has been renamed to C/S Voyager. Operational
R/V milliAmpere1 [2] Operational
R/V milliAmpere2 [5] Operational
R/V Gunnerus [17] Operational
Table 1: The C/S Fleet and research vessels at the time of writing with relevant references.
To complement these physical assets, we have developed corresponding simulation environments for each vessel. These digital platforms support early-stage algorithm development and debugging, ranging from simple integration-test models to high-fidelity simulations with realistic hydrodynamic effects. Additionally, sensor-model simulations enable testing of algorithms that rely on complex sensor feedback. We also employ a digital twin environment in Unity, allowing us to visualize and analyze vessel behaviors and gain better insights.
Taken together, these components form our design, verification, test, and validation pipeline. Algorithms are first iterated upon in simulation, where rapid testing reduces cost and risk, before being transferred to the physical model vessels in our laboratory basin. This progressive approach not only streamlines the overall testing process but also ensures that MASS algorithms are thoroughly verified and validated across increasing levels of realism and operational complexity prior to deployment in real maritime conditions.
3 Digital infrastructure
3.1 Simulation environment
To support algorithm development and early-stage testing, our infrastructure includes three complementary simulation environments, each tailored for different stages of the development and verification process. These simulation environments serve distinct purposes but are all aligned within our design and verification pipeline, ensuring that algorithms can be tested efficiently from conceptual stages to realistic operational scenarios before transitioning to physical trials.
3.1.1 High-fidelity simulation model
The high-fidelity simulation framework mcsimpy [^2] The mcsimpy open-source can be accessed at https://doi.org/10.5281/zenodo.17274093, written in Python, replicates the dynamics of the C/S fleet by incorporating wave-excitation models and realistic hydrodynamic coefficients. Specifically, it uses JONSWAP [12] and Pierson–Moskowitz [25] wave spectra to generate diverse sea states, while hydrodynamic parameters are derived from WAMIT [36] and VERES [30]. mcsimpy supports both 3 and 6 degrees of freedom (DOF) models, making it suitable for various vessel configurations, including the R/V Gunnerus. Built on the studies by [19, 14] as well as the Marine Systems Simulator [23], this framework provides a versatile platform for simulating control and navigation algorithms under realistic environmental conditions. The intended use of the simulation environment is for control algorithm development, controller tuning, and serving as a fast, flexible sandbox for machine learning (ML) and reinforcement learning (RL) applications.
3.1.2 Reduced-order model simulation
We employ a lightweight simulation environment [^3] The simulator for the simplified hydrodynamic model is open-source and available at https://doi.org/10.5281/zenodo.17277252 that approximates vessel dynamics using reduced-order models. This environment is primarily used for rapid testing of software interfaces and high-level control logic, as it shares the same programming interface as the physical platform.
Each vessel is treated as a uniform rectangular prism of dimensions $L$, $B$, and $T$, giving a mass of $m=\rho LBT$, with principal moments of inertia:
$I_{x}=\tfrac{1}{12}m(B^{2}+T^{2}),\quad I_{y}=\tfrac{1}{12}m(L^{2}+T^{2}),\quad I_{z}=\tfrac{1}{12}m(L^{2}+B^{2}).$ (1)
These form the diagonal rigid-body mass–inertia matrix $\mathbf{M}_{\mathrm{RB}}$. To approximate hydrodynamic effects, diagonal added-mass terms $\mathbf{M}_{\mathrm{A}}$ and linear damping coefficients $\mathbf{D}$ are included. The added-mass and damping terms are estimated by scaling the corresponding rigid-body values by predefined factors. For example, the added mass in surge is typically approximated as 20% of the rigid-body mass, while in sway and roll, it is often set to approximately 100% of the rigid-body mass. Thruster dynamics are omitted for computational efficiency. The model employs the standard 6DOF nonlinear equations of motion,
$\dot{\eta}=\mathbf{J}(\eta)\nu,$ (2)
$(\mathbf{M}_{\mathrm{RB}}+\mathbf{M}_{\mathrm{A}})\dot{\nu}+\bigl(\mathbf{C}_{\mathrm{RB}}(\nu)+\mathbf{C}_{\mathrm{A}}(\nu_{r})\bigr)\nu_{r}+\mathbf{D}\nu_{r}=\tau+\mathbf{g}(\eta),$ (3)
where the Coriolis matrices $\mathbf{C}_{\mathrm{RB}}$ and $\mathbf{C}_{\mathrm{A}}$ incorporate rigid-body and added-mass terms, respectively, $\tau$ represents loads acting on the body, and $\mathbf{J}$ is the transformation matrix between the body and inertial frames [8]. For small roll and pitch angles, a diagonal restoring term $\mathbf{g}(\eta)$ provides simple restoring forces in heave, roll, and pitch. Numerically, we integrate $\eta$ and $\nu$ using a fourth-order Runge–Kutta step at each timestep $\Delta t$. Although this reduced-order approach omits detailed thruster and environmental modeling, it offers sufficient fidelity for early-stage verification and greatly reduces computational overhead.
3.1.3 Simplified dynamics with sensor models
Stonefish [3] is a high performance and real-time capable robotics simulator written in C++. Here, it is used for providing a sandbox for image processing, supports a wide range of sensor types (such as cameras, inertial measurement units, and sonars), and has simplified collision and wave dynamics. Thanks to its flexibility, Stonefish is ideal for evaluating algorithms like localization, object detection, and platform interaction that requires environmental variability and realistic sensor feedback. Stonefish bridges the gap between reduced-order simulation models and large-scale real-world tests by simulating important elements of the environment, and allows validation of control and perception algorithms.
Figure 2: C/S Voyager (on the left) and C/S Enterprise I (on the right) in the Stonefish simulation environment.
3.2 Digital Twin
A digital twin environment has been developed to support testing and visualization of vessel behavior in a virtual representation of the MC-Lab  [26]. It consists of Unity-implemented 3D models of the MC-Lab and selected vessels from the C/S fleet, including the C/S Enterprise I, C/S Saucer, and C/S Inocean Cat I Drillship. This setup enables the monitoring and demonstration of control algorithms in a more intuitive format and can be used to support testing in early development phases.
The digital twin system is paired with a remote control center, which allows operators to control the physical vessels in the MC-Lab from a separate location [11]. Command can be performed manually using joysticks or Metaquest 2 headset and hand controls [^4] Link to the video demonstration: https://www.youtube.com/watch?v=fqEdp8VXAqE, as well as using these interfaces to provide guidance commands like waypoints and speed references for automatic control. Vessel state information, such as position and orientation, is transmitted over the site-to-site network to provide remote monitoring capabilities. While some limitations exist due to hardware constraints—particularly in the responsiveness of the monitoring—the system has been tested in both simulation and physical experiments and has been shown to function reliably under the available conditions.
3.3 Software
3.3.1 Core modules
Figure 3: Control architecture of the testbed environment showing the closed-loop system: sensors feed an EKF observer that drives controllers, whose outputs route through thrust allocation to actuators. A multiplexer allows switching between control modules.
All the vessels in the C/S fleet operate on a shared software stack [^5] C/S software suite can be accessed at https://doi.org/10.5281/zenodo.17233653. The software is developed in a monolithic architecture, with all components maintained in a single repository. This structure simplifies version control, integration, and testing, and is supported by a continuous integration and deployment (CI/CD) pipeline that automates the build and deployment process for each vessel in the testbed.
The core software is written in Python and C++, combining the flexibility and rapid development capabilities of Python with the performance and low-level control afforded by C++. Robot Operating System 2  [16] (ROS 2) serves as the middleware layer, enabling efficient inter-process communication and seamless integration with additional components and external programs. This middleware facilitates modularity, allowing various control, sensing, and estimation modules to operate independently while maintaining synchronized communication across the system.
3.3.2 Control and estimation modules
Each vessel comes with a set of default software modules, allowing for immediate testing and validation of standard behaviors. These include a thrust allocation module for distribution of forces to the thrusters, a single-input-single-output (SISO) proportional-integrated-derivative (PID) based feedback velocity controller, third-order reference filter integrated with a proportional-derivative (PD) pose controller, and finally an extended Kalman filter (EKF) from the robot_localization package [20].
The C/S fleet shares a unified and modular control architecture that simplifies integration into the broader testbed infrastructure and provides a consistent foundation for algorithm development. Built-in control and estimation modules offer baseline performance metrics, serving as reference implementations for testing and benchmarking new strategies. Depending on the experimental need, these modules can be disabled or replaced with custom implementations, allowing researchers to test their own algorithms in a consistent environment. To this end, we choose to keep the default modules as simple as possible, while still providing the necessary functionality for basic control and estimation tasks. The overall control architecture is illustrated in Fig. 3. Using this architecture, we can easily test and validate various control algorithms consistently.
Thrust allocation
We employ the maneuvering-based dynamic thrust allocation method [9], which formulates the thrust allocation task as a dynamic control problem. This approach uses a control Lyapunov function (CLF) and a control barrier function (CBF) to handle rate limitations and force saturations, ensuring smooth and feasible thrust commands.
Velocity controller
A proportional-integral with reference feedforward (PI-RFF) velocity controller is made available to track a desired velocity. For the body-frame velocity $\nu:=[u,v,r]^{\top}\in\mathbb{R}^{3}$, the 3-DOF surge–sway–yaw plant without Coriolis or centrifugal terms is
$\mathbf{M}\dot{\nu}+\mathbf{D}\nu=\tau,$ (4)
where $\mathbf{M}=\mathbf{M}^{\top}>0$ and $\mathbf{D}>0$. Let the reference velocity be $\nu_{d}(t)$, define the tracking error $\tilde{\nu}:=\nu-\nu_{d}$, the integral state $\dot{\xi}=\tilde{\nu}$, and choose $a$ as the virtual control input,
$a:=\dot{\nu}_{d}-K_{p}\tilde{\nu}-K_{i}\xi,$ (5)
where $K_{p}=\operatorname{diag}(k_{p1},k_{p2},k_{p3})>0$ and $K_{i}=\operatorname{diag}(k_{i1},k_{i2},k_{i3})>0$. The commanded loads (forces/torques) become
$\tau_{\mathrm{cmd}}=\mathbf{M}a+\mathbf{D}\nu_{d},$ (6)
which combines acceleration reference feedforward $\mathbf{M}\dot{\nu}_{d}$, proportional-integral feedback, and feedforward compansation of damping using $\mathbf{D}\nu_{d}$. Consider the Lyapunov function,
$V(\tilde{\nu},\xi)=\tfrac{1}{2}\tilde{\nu}^{\top}\mathbf{M}\tilde{\nu}+\tfrac{1}{2}\xi^{\top}K_{i}\mathbf{M}\xi,$ (7)
and its derivative
$\dot{V}(\tilde{\nu},\xi)=-\tilde{\nu}^{\top}(\mathbf{D}+\mathbf{M}K_{p})\tilde{\nu}\leq 0.$ (8)
Because $\mathbf{D}+\mathbf{M}K_{p}>0$, the Lyapunov function $V$ is positive-definite and radially unbounded, while $\dot{V}$ is negative semi-definite. A straightforward application of the Krasovskii-LaSalle invariance principle will show that the equilibrium $(\tilde{\nu},\xi)=(0,0)$ is uniformly globally asymptotically stable.
Pose control
We employ a mass-damper-spring based reference model for the stationkeeping controller. The reference filter defines a state vector $\mathbf{x}(t)=[\eta^{\top}_{d}(t),\dot{\eta}^{\top}_{d}(t),\ddot{\eta}^{\top}_{d}(t)]^{\top}$, where $\eta_{d}\in\mathbb{R}^{3}$ is the desired pose in the world frame. The filter evolves according to the continuous-time state-space model
$\dot{\mathbf{x}}(t)=\mathbf{A}_{d}\mathbf{x}(t)+\mathbf{B}_{d}\eta_{r},$ (9)
where $\eta_{r}(t)$ is the desired steady-state pose input. The filter matrices are defined as
$\mathbf{A}_{d}=\begin{bmatrix}\mathbf{0}_{3\times 3}&\mathbf{I}_{3}&\mathbf{0}_{3\times 3}\\ \mathbf{0}_{3\times 3}&\mathbf{0}_{3\times 3}&\mathbf{I}_{3}\\ -\mathbf{\Omega}^{3}&-\bigl(2\mathbf{\Delta}+\mathbf{I}_{3}\bigr)\mathbf{\Omega}^{2}&-\bigl(2\mathbf{\Delta}+\mathbf{I}_{3}\bigr)\mathbf{\Omega}\end{bmatrix},\qquad\mathbf{B}_{d}=\begin{bmatrix}\mathbf{0}_{3\times 3}\\ \mathbf{0}_{3\times 3}\\ \mathbf{\Omega}^{3}\end{bmatrix}$, (10)
with $\mathbf{\Omega}=\mathrm{diag}(\omega_{1},\omega_{2},\omega_{3})$ controlling the filter bandwidth, and $\mathbf{\Delta}=\mathrm{diag}(\delta_{1},\delta_{2},\delta_{3})$ controlling the filter damping [8]. Other types of reference models may also be implemented; see [7]. A forward-Euler discretization step is sufficient to update the filter state, providing smooth desired pose, velocity, and acceleration outputs for the DP controller. The default included controller is simply a PD control law to generate the desired thrust loads for the vessel to closely follow the reference trajectory using
$\tau_{cmd}=-\mathbf{K}_{p}\mathbf{R}(\psi)^{\top}\bigl(\eta-\eta_{d}\bigr)-\mathbf{K}_{d}\mathbf{R}(\psi)^{\top}\bigl(\dot{\eta}-\dot{\eta}_{d}\bigr)$ (11)
$\mathbf{R}(\psi)=\begin{bmatrix}\cos\psi&-\sin\psi&0\\ \sin\psi&\cos\psi&0\\ 0&0&1\end{bmatrix}$, (12)
where $\tau_{cmd}\in\mathbb{R}^{3}$ is the thrust load control vector, and $\mathbf{K}_{p}=\mathbf{K}_{p}^{\top}>0$ and $\mathbf{K}_{d}=\mathbf{K}_{d}^{\top}>0$ are the proportional and derivative gain matrices, respectively. The results of the pose controller are shown in Fig. 4 for both the reduced-order simulation and physical model in the basin.
(a) Reduced-order simulation
(b) Basin experiment
Figure 4: Four corner tests for C/S Voyager in simulation (left) and physical basin (right).
The 4-corner test [33] was conducted using the pose controller, and the results in Fig. 4 show the performance differences between simulation and physical implementation. The simulation achieved a position RMSE of 0.062m, yaw RMSE of $0.439^{\circ}$, and velocity RMSE of 0.031m/s, whereas the physical vessel showed 0.180m, $4.295^{\circ}$, and 0.036m/s, respectively.
Controller selection
A multiplexer is used to choose between different control modes. This enables operators to choose the most appropriate control strategy for a given task on the fly.
3.4 Available data
To build realistic simulators of the C/S fleet, we have generated and publicly released comprehensive hydrodynamic datasets [^6] Dataset is accessible through https://doi.org/10.5281/zenodo.17274087. We performed stationkeeping hydrodynamic analyses using WAMIT [36]; a boundary-element software grounded in potential flow theory. The submerged portions of each hull were meshed from approximate 3D scans with some geometric simplifications. To optimize computation, we simulated only half of each hull by exploiting ship symmetry. From these runs, we obtained hydrostatic coefficients, linear hydrodynamic loads, second-order mean drift forces and moments, and response amplitude operators (RAOs) for the six rigid-body motions. Linear hydrodynamic loads (added mass and damping) were derived from the radiation problem, while wave excitations were computed from the diffraction problem over headings ranging from $0^{\circ}$ to $180^{\circ}$. The RAOs were based on vessel mass alone, excluding mooring or viscous effects.
4 Physical infrastructure
The MC-Lab provides a versatile and well-instrumented environment for testing model-scale marine vessels in realistic wave conditions. Its primary facility is a wave basin measuring $32m$ from the flap-type wave maker to the beach and 6.4m wide, with a maximum depth of $1.5m$. The wave maker flap spans the full basin width, and is capable of generating waves up to $0.25m$ in amplitude at controlled frequencies, and a wave-damping beach is located at the far end that minimizes reflections for cleaner experiments. This allows users to produce both regular and irregular wave spectra, such as JONSWAP, in either manual or automated modes. Conductive and ultrasonic probes can be placed throughout the basin to monitor wave characteristics. A Motion Capture system offers real-time object tracking (to within $1cm$ in position and $0.5^{\circ}$ in orientation). Additional technical specifications and usage details can be found on the web pages of [4]. Within 2030, the MC-Lab will be succeeded by a new model basin for academic research, called the Hydro-Cybernetics Laboratory in the House of Archimedes as part of the new Norwegian Ocean Technology Centre [22]. This more modern facility will be a larger test basin, approximately $65m$ in length, $8m$ in width, and $3m$ in depth, further expanding research capabilities in hydrodynamics, marine cybernetics, and ocean technology.
4.1 Hardware
All vessels in the C/S fleet share a standardized hardware architecture designed for modularity, scalability, and ease of integration. The primary differences between the vessels are the number and types of actuators and sensors, which are abstracted through a common software stack. This allows all vessels to operate with minimal configuration differences, streamlining development and deployment across the fleet. As a result, new vessels can be added to the testbed environment with minimal re-engineering, using the existing architecture as a template for both hardware and software. This consistency reduces development time and increases maintainability, reliability, and scalability.
At the core of each vessel is a Raspberry Pi single-board computer, which serves as the main onboard processor. This choice offers a compact, capable, and well documented platform for running control algorithms, handling sensor input, and managing actuator commands in real time. Communication between onboard components is kept streamlined and maintainable by using standard hardware protocols, ensuring compatibility with a wide range of off-the-shelf sensors and devices; see Fig. 5.
Figure 5: Shared hardware architecture across the C/S fleet.
5 Conclusions
In this paper, we presented our development of a digital-physical testbed for ship autonomy research. The testbed is designed to support the full pipeline of algorithm development, verification, and validation for MASS in a controlled and flexible environment. The infrastructure consists of a fleet of small-scale model vessels, vessel-specific simulation environments, and a digital twin implemented in Unity. These components are tightly integrated into a modular and scalable testbed, allowing researchers to transition seamlessly between simulation and physical experimentation. The testbed enables safe, repeatable, and cost-effective testing of autonomy algorithms, while also supporting more advanced use cases such as remote control and hybrid testing scenarios.
While the digital-physical testbed offers a valuable platform for development and testing, it has certain limitations. First, small-scale models do not fully capture the behavior of full-scale systems, and simulations can only approximate real-world conditions. Moreover, a bollard-pull test—which is commonly used in marine engineering to measure the thrust generated by a vessel’s propulsion system at rest—has not yet been conducted, limiting the accuracy of the thrust modeling. The controlled laboratory environment also restricts exposure to the variability and unpredictability of real maritime operations. Lastly, the testbed’s resource requirements may pose challenges to its scalability.
We plan to expand the simulation capabilities by incorporating forward-speed hydrodynamic data from ShipX [30], extending high-fidelity simulators to cover relevant speed ranges for the C/S fleet, and publishing experimental datasets for benchmarking and reproducibility.
6 Data Availability Statement
The datasets and software that support the findings of this study are openly available. The hydrodynamic datasets used for model development can be accessed at https://doi.org/10.5281/zenodo.17274087. The high-fidelity simulation framework mcsimpy is available at https://doi.org/10.5281/zenodo.17274093, the reduced-order simulator shoeboxpy at https://doi.org/10.5281/zenodo.17277252, and the C/S software suite at https://doi.org/10.5281/zenodo.17233653. All repositories are publicly accessible under open-source licenses.
7 Acknowledgments
Work funded by the Research Council of Norway (RCN) through SFI AutoShip(RCN project 309230), NTNU AMOS (RCN project 223254), SFI Harvest (RCN project 309661), and partly funded by the European Union via the LORELEI-X Twinning project (GA 101159489) and NTNU VISTA CAROS. Authors would also like to thank Jon Estil Krågebakk for providing 3D models, and the team of engineers at MC-Lab, Robert Opland, and Vebjørn Steinsholt.